# Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit

### v2

@@
Methods to achieve near-millisecond energy relaxation and dephasing times for a superconducting transmon qubit
Mikko Tuokkola Corresponding author email: mikko.tuokkola@aalto.fi    Yoshiki Sunada    Heidi Kivijärvi    Jonatan Albanese QCD Labs, QTF Centre of Excellence, Department of Applied Physics,
Aalto University, P.O. Box 13500, FIN-00076 Aalto, Finland
   Leif Grönberg    Jukka-Pekka Kaikkonen    Visa Vesterinen    Joonas Govenius VTT Technical Research Centre of Finland Ltd. & QTF Centre of Excellence, P.O. Box 1000, 02044 VTT, Finland    Mikko Möttönen Corresponding author email: mikko.mottonen@aalto.fi QCD Labs, QTF Centre of Excellence, Department of Applied Physics,
Aalto University, P.O. Box 13500, FIN-00076 Aalto, Finland
VTT Technical Research Centre of Finland Ltd. & QTF Centre of Excellence, P.O. Box 1000, 02044 VTT, Finland
(13. syyskuuta 2024)
Tiivistelmä
Superconducting qubits are one of the most promising physical systems for implementing quantum computers. However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits. Here, we report our measurement results of a high-coherence transmon qubit with energy relaxation and echo dephasing times surpassing those in the existing literature. We measure a qubit frequency of 2.890 GHz, an energy relaxation time $T_{1}$ with a median of $502$ µs and a maximum of $(765\pm 82.6)$ µs, and an echo dephasing time $T_{2}^{\mathrm{echo}}$ with a median of $541$ µs and a maximum of $(1057\pm 138)$ µs. We report in detail our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
@@ I Introduction
Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations. The longest reported energy relaxation times $T_{1}$ of transmons have approached but not surpassed 500-µs median value as shown in Fig. 1 and Extended Data Table 1 based on Refs. [9, 10, 11, 12, 13, 14, 15, 16]. Recently, there have also been significant improvements in transmon echo dephasing times $T_{2}^{\mathrm{echo}}$ with the longest reported average extending to 307 µs [10, 13, 14, 16]. Transmon qubits with coherence times above 100 µs have also been reported, for example, in Refs.  [17, 18, 19, 20, 21, 22]. With other types of superconducting qubits such as 0-$\pi$ qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [23, 24, 25].
Kuva 1: Summary of the longest measured energy relaxation and echo dephasing times in the existing literature. Filled squares represent the median or average values, and empty squares represent the longest values reported in the references which are chronological ordered.
@@ II.0.1 Substrate and niobium patterning
The Nb film is patterned in a plasma processing system by a chemical dry-etching process, clearing the areas revealed by the resist development step. Immediately before the process, the empty plasma chamber is cleaned with a combination of CF4 and O2 gases for a total of 5 min. We then load the sample into the chamber and apply oxygen plasma ashing for 10 s at 100 mTorr chamber pressure, 40 sccm gas flow, and 150 W rf source power to remove the post-development residue of photoresist on the sample surface. Then, we pump out the oxygen and introduce CF4 at 50-mTorr chamber pressure, 20-sccm flow, and 30-W source power. To ensure that the Nb film is fully etched, we carry out an additional etch after unloading, visually inspecting, and re-loading the sample. The etch rate of Si is significantly lower than that of Nb, which helps us to achieve a convenient control of the etch depth. After the etching process, we remove some of the residual chemicals by applying another oxygen plasma ashing for 2 min without breaking the vacuum.
@@ II.0.3 Junction deposition
For the bottom layer of the junctions, we deposit 40 nm of Al at $\theta=45^{\circ}$ tilt and $\phi=-45^{\circ}$ planetary angle. Then we create the insulating aluminum oxide layer by static oxidation at 1.2 mbar for 5 min. After the oxidation, we deposit the second Al layer in two steps with $\theta=45^{\circ}$ tilt and two planetary angles: first 30 nm at $\phi=45^{\circ}$ and then 30 nm at $\phi=-135^{\circ}$. This ensures that both sides of the oxidized bottom strip are covered by the second Al layer.
@@ III Setup for qubit measurements
Our experimental setup used for qubit characterization and measurements is presented in Fig. 3, with detailed information on measurement equipment and components provided in Extended Data Table 3. The sample is wire-bonded to a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines). The sample is exposed to ambient atmosphere at room temperature for a total of 7 days. It is then cooled down to approximately 10 mK in a Bluefors dilution refrigerator equipped with a tin-plated copper shield at 10 mK and an Amumetal magnetic shield just inside the outer vacuum chamber (OVC). We directly generate the qubit control and readout signals without any analog mixers by using a Xilinx RFSoC evaluation board with QICK firmware [33, 34] locked to a Rb frequency standard. The signals then pass through rf attenuators, an Eccosorb infrared filter, and a low pass filter before reaching the sample. The readout signal coming out of the sample goes through a three-wave-mixing Josephson traveling-wave parametric amplifier (TWPA), a cryogenic high-electron-mobility-transistor (HEMT) amplifier, and a room-temperature HEMT amplifier before being digitized by RFSoC. During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
Kuva 3: Schematic of the experimental setup for qubit characterization and measurement. See Extended Data Table 3 for a description of each component. For the second cooldown, the $10$-dB attenuator in the drive line at the mixing-chamber (MXC) plate is removed, and the first $10$-dB attenuator in the TWPA pump line at the MXC plate is replaced with an Eccosorb filter.
Taulukko 1: Summary of the measurement results for all four qubits during the first cooldown. Here, $T_{1}$ and $T_{2}^{\mathrm{echo}}$ are median values. The short $T_{1}$ and $T_{2}^{\mathrm{echo}}$ of qubit Q3 can be explained by the fact that the aluminum film evaporated on top of the resist mask inside its SQUID loop is not removed by the lift-off process.
@@
$\mathrm{Q}_{2}$ (fixed) 2.890 502 549 9.1
$\mathrm{Q}_{3}$ (tunable) 4.593 44 28 1.3
@@
(a)(b)(e)(f)(c)(d)(g)(h)
Kuva 4: Energy relaxation time and echo dephasing time of qubit $\bm{\mathrm{Q}_{2}}$. (a–d) Distributions of (a, c) $T_{1}$ and (b, d) $T_{2}^{\textrm{echo}}$ during the (a, b) first and (c, d) second cooldown. (e, g) Time trace for the longest measured energy relaxation time during the (e) first and (g) second cooldown resulting in $T_{1}=(765\pm 83)$ µs and $(666\pm 33)$ µs, respectively. (f, h) Time trace for the longest measured echo dephasing time during the (f) first and (h) second cooldown resulting in $T_{2}^{\mathrm{echo}}=(1057\pm 138)$ µs and (806±78(806$\pm$ 78( 806 ± 78) µs, respectively. Panels (c) and (d) include results combined with and without TWPA pump and panel (g) has been obtained with the pump. For other panels, the pump has not been applied.
After the first cooldown, we reconfigure the attenuators as described in the caption of Fig. 3 before starting another cooldown. During the second cooldown, we compare measurements with and without a pump signal applied to the TWPA. The 10.5-GHz pump signal passes through a 4–12 GHz circulator, reflects from a 8-GHz lowpass filter, and travels through the circulator again before reaching the TWPA. This configuration allows us to combine the pump and readout signals without the power dissipation of a directional coupler or the impedance mismatch of a diplexer.
@@ IV Measurement Results
We measure the fundamental modes of the four readout resonators to have frequencies ranging from $5.85$ GHz to $6.23$ GHz and the fixed-frequency qubits to have qubit frequencies of $f_{\mathrm{Q}_{2}}=2.89$ GHz and $f_{\mathrm{Q}_{4}}=3.29$ GHz. The flux-tunable qubits have maximum frequencies of $f_{\mathrm{Q}_{1}}=4.05$ GHz and $f_{\mathrm{Q}_{3}}=4.59$ GHz which we find by sweeping the external magnetic flux applied to the sample. We then follow the standard procedure for measuring the energy relaxation time $T_{1}$ and echo dephasing time $T_{2}^{\mathrm{echo}}$ of each qubit [2].
During the first cooldown, we measure that the fixed-frequency qubit $Q_{2}$ exhibits especially long energy relaxation and echo dephasing times: medians of $T_{1}=502$ µs and $T_{2}^{\mathrm{echo}}$ of $541$ µs and maxima of $T_{1}=(765\pm 0.083)$ µs and $T_{2}^{\mathrm{echo}}=(1057\pm 0.138)$ µs. The data for these results are shown in Fig. 4(a, b, e, f) in the form of distributions for repeated measurements of $T_{1}$ and $T_{2}^{\mathrm{echo}}$ and time traces for the maximum values. In each $T_{1}$ time trace, we repeat and average 500 measurements for each value of delay time $T_{\mathrm{delay}}$ between a $\pi$-pulse and a readout pulse. The number of repetitions is $2000$ for the $T_{2}^{\mathrm{echo}}$ measurement. The measurement results for all four qubits during the first cooldown is summarized in Table 1.
@@
The data for the coherence time measurements carried out in this section are available in Ref. [28].
@@ V Discussion
In this work, we presented detailed information on our design, fabrication methods, and measurement setup for a high-coherence transmon qubit. One of the four qubits on the measured sample achieved a median energy relaxation time $T_{1}$ of $502$ µs and a median echo dephasing time $T_{2}^{\textrm{echo}}$ of $541$ µs. These results surpass the previous results for a transmon qubit reported in the literature. However, the $T_{1}$ and $T_{2}^{\mathrm{echo}}$ were significantly shorter in the second cooldown. This may be due to the redistribution of environmental fluctuators and oxidation of the sample surface. Our fabrication method and experimental setup can also be applied to other types of superconducting qubits such as the unimon qubit to enhance its energy relaxation and dephasing times [35].
In conclusion, this result represents a significant step in the development of high-coherence superconducting qubits by approaching the millisecond mark for the energy relaxation and dephasing times of a transmon qubit. It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref. 16 and the effectiveness of the equipment and components listed in Extended Data Tables 2 and 3. Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.

### v3

@@
Abstract
Superconducting qubits are one of the most promising physical systems for implementing quantum computers. However, executing quantum algorithms of practical computational advantage requires further improvements in the fidelities of qubit operations, which are currently limited by the energy relaxation and dephasing times of the qubits. Here, we report our measurement results of a high-coherence transmon qubit with energy relaxation and echo dephasing times surpassing those in the existing literature. We measure a qubit frequency of 2.9 GHz, an energy relaxation time $T_{1}$ with a median of $425$ µs and a maximum of $(666\pm 33)$ µs, and an echo dephasing time $T_{2}^{\mathrm{echo}}$ with a median of $541$ µs and a maximum of $(1057\pm 138)$ µs. We report in detail our design, fabrication process, and measurement setup to facilitate the reproduction and wide adoption of high-coherence transmon qubits in the academia and industry.
@@ I Introduction
Short coherence times have historically been a disadvantage of superconducting qubits, but the previous decade has shown continuous improvements on this challenge, leading to increased fidelities of qubit operations. The longest reported energy relaxation times $T_{1}$ of transmons have approached but not surpassed 400-µs median value as shown in Fig. 1 and Extended Data Table 1 based on Refs. [25, 34, 11, 7, 28, 4, 2, 18]. Recently, there have also been significant improvements in transmon echo dephasing times $T_{2}^{\mathrm{echo}}$ with the longest reported average extending to 307 µs [34, 28, 4, 18]. Transmon qubits with coherence times above 100 µs have also been reported, for example, in Refs.  [9, 35, 22, 36, 21, 24]. With other types of superconducting qubits such as 0-$\pi$ qubits and fluxonium qubits, energy relaxation times above a millisecond have been observed, but these qubits have either had short dephasing times or low qubit frequencies [26, 12, 29].
Figure 1: Summary of the longest measured energy relaxation and echo dephasing times in the existing literature. Squares represent the median or average values, and stars represent the longest values reported in the references which are chronologically ordered.
@@ II.0.1 Substrate and niobium patterning
The Nb film is patterned in a plasma processing system by a chemical dry-etching process, clearing the areas revealed by the resist development step. Immediately before the process, the empty plasma chamber is cleaned with a combination of CF4 and O2 gases for a total of 5 min. We then load the sample into the chamber and apply oxygen plasma ashing for 10 s at 100 mTorr chamber pressure, 40 sccm gas flow, and 150 W rf source power to remove the post-development residue of photoresist on the sample surface. Then, we pump out the oxygen and introduce CF4 at 50-mTorr chamber pressure, 20-sccm flow, and 30-W source power. To ensure that the Nb film is fully etched, we carry out an additional etch after unloading, visually inspecting, and re-loading the sample. The etch rate of Si is significantly lower than that of Nb, which helps us to achieve a convenient control of the etch depth. The typical total etching time of Nb film is around 10 min. After the etching process, we remove some of the residual chemicals by applying another oxygen plasma ashing for 2 min without breaking the vacuum.
@@ II.0.3 Junction deposition
For the bottom layer of the junctions, we deposit 40 nm of Al at $\theta=45^{\circ}$ tilt and $\phi=-45^{\circ}$ planetary angle. Then we create the insulating aluminum oxide layer by static oxidation at 1.2 mbar for 5 min. During the oxidation, the aluminum source is protected from the oxidation by a valve that blocks the oxygen flow. After the oxidation, we deposit the second Al layer in two steps with $\theta=45^{\circ}$ tilt and two planetary angles: first 30 nm at $\phi=45^{\circ}$ and then 30 nm at $\phi=-135^{\circ}$. This ensures that both sides of the oxidized bottom strip are covered by the second Al layer.
@@ III Setup for qubit measurements
Our experimental setup used for qubit characterization and measurements is presented in Fig. 3, with detailed information on measurement equipment and components provided in Extended Data Table 3. The sample is wire-bonded to the printed circuit board (PCB) of a QCage.24 sample holder and placed inside QCage Magnetic Shielding, both of which are supplied by QDevil (under Quantum Machines). The chip is suspended by four corners inside a cavity and clamped down by the PCB. The assembly is placed inside a light-tight superconducting aluminium enclosure to reject stray microwave and infrared photons before being mounted inside the magnetic shield. The sample is exposed to ambient atmosphere at room temperature for a total of 7 days. It is then cooled down to approximately 10 mK in a Bluefors dilution refrigerator equipped with a tin-plated copper shield at 10 mK and an Amumetal magnetic shield just inside the outer vacuum chamber (OVC). We directly generate the qubit control and readout signals without any analog mixers by using a Xilinx RFSoC evaluation board with QICK firmware [30, 8] locked to a Rb frequency standard. The signals then pass through rf attenuators, an Eccosorb infrared filter, and a low-pass filter before reaching the sample. The readout signal coming out of the sample goes through a three-wave-mixing Josephson traveling-wave parametric amplifier (TWPA), a cryogenic high-electron-mobility-transistor (HEMT) amplifier, and a room-temperature HEMT amplifier before being digitized by RFSoC. During the first cooldown, the TWPA is not pumped and therefore does not provide any amplification.
Figure 3: Schematic of the experimental setup for qubit characterization and measurement. See Extended Data Table 3 for a description of each component. For the second cooldown, the $10$-dB attenuator in the drive line at the mixing-chamber (MXC) plate is removed, and the first $10$-dB attenuator in the TWPA pump line at the MXC plate is replaced with an Eccosorb filter.
Table 1: Summary of the measurement results for all four qubits during the first cooldown. Here, $T_{1}$ and $T_{2}^{\mathrm{echo}}$ are median values. The short $T_{1}$ and $T_{2}^{\mathrm{echo}}$ of qubit Q3 can be explained by the fact that the aluminum film evaporated on top of the resist mask inside its SQUID loop is not removed by the lift-off process.
@@
$\mathrm{Q}_{2}$ (fixed) 2.890 N/A [^1] During the first cooldown, $T_{1}$ of the qubit $\mathrm{Q}_{2}$ was measured with such a short waiting time that the qubit did not decay entirely before the next measurement. 549 9.1
$\mathrm{Q}_{3}$ (tunable) 4.593 44 32 1.3
@@
(a)(b)(d)(e)(c)(f)
Figure 4: Energy relaxation time and echo dephasing time of qubit $\bm{\mathrm{Q}_{2}}$. (a–c) Distributions of (b) $T_{1}$ and (a, c) $T_{2}^{\textrm{echo}}$ during the (a) first and (b, c) second cooldown. (e) Time trace for the longest measured energy relaxation time during the (e) second cooldown resulting in $T_{1}=(666\pm 33)$ µs. (d, f) Time trace for the longest measured echo dephasing time during the (d) first and (f) second cooldown resulting in $T_{2}^{\mathrm{echo}}=(1057\pm 138)$ µs and (806±78(806$\pm$ 78( 806 ± 78) µs, respectively. Panels (b) and (c) include results combined with and without TWPA pump and panel (e) has been obtained with the pump. For other panels, the pump has not been applied.
After the first cooldown, we reconfigure the attenuators as described in the caption of Fig. 3 before starting another cooldown. During the second cooldown, we compare measurements with and without a pump signal applied to the TWPA. The 10.5-GHz pump signal passes through a 4–12 GHz circulator, reflects from a 8-GHz low-pass filter, and travels through the circulator again before reaching the TWPA. This configuration allows us to combine the pump and readout signals without the power dissipation of a directional coupler or the impedance mismatch of a diplexer.
@@ IV Measurement Results
We measure the fundamental modes of the four readout resonators to have frequencies ranging from $5.85$ GHz to $6.23$ GHz and the fixed-frequency qubits to have qubit frequencies of $f_{\mathrm{Q}_{2}}=2.89$ GHz and $f_{\mathrm{Q}_{4}}=3.29$ GHz. The flux-tunable qubits have maximum frequencies of $f_{\mathrm{Q}_{1}}=4.05$ GHz and $f_{\mathrm{Q}_{3}}=4.59$ GHz which we find by sweeping the external magnetic flux applied to the sample. We then follow the standard procedure for measuring the energy relaxation time $T_{1}$ and echo dephasing time $T_{2}^{\mathrm{echo}}$ of each qubit [19]. The flux-tunable qubits are measured in the qubit sweet-spot, and the frequency dependence of the relaxation and dephasing times is left for future work.
[…]
@@
The data for the coherence time measurements carried out in this section are available in Ref. [31]. In the Supplementary Note S1 of the Supplementary Information, we demonstrate the reproducibility of the high-coherence fabrication process by measuring additional transmon qubits in three-dimensional cavity resonators.
@@ V Discussion
In this work, we presented detailed information on our design, fabrication methods, and measurement setup for a high-coherence transmon qubit. One of the four qubits on the measured sample achieved a median energy relaxation time $T_{1}$ of $425$ µs and a median echo dephasing time $T_{2}^{\textrm{echo}}$ of $541$ µs. These results surpass the previous results for a transmon qubit reported in the literature. However, the $T_{1}$ and $T_{2}^{\mathrm{echo}}$ were significantly shorter in the second cooldown. This may be due to the redistribution of environmental fluctuators and oxidation of the sample surface. Our fabrication method and experimental setup can also be applied to other types of superconducting qubits, such as the unimon qubit, to enhance its energy relaxation and dephasing times [14], and also to large-scale manufacturing as in Ref. [32].
In conclusion, this result represents a significant step in the development of high-coherence superconducting qubits by approaching the millisecond mark for the energy relaxation and dephasing times of a transmon qubit. It also demonstrates the robustness and reproducibility of the fabrication recipe described in Ref. 18 and the effectiveness of the equipment and components listed in Extended Data Tables 2 and 3. Detailed reporting of a high-coherence qubit will benefit the research community and accelerate the global efforts on developing quantum sensors, quantum simulators, and quantum computers based on superconducting quantum technologies.